Is y = (x + 3)(x – 6) – 2x2 linear or quadratic? Identify the quadratic, linear, and constant terms.
Answer by richwmiller(17219)

quadratic equations have squared variables(x's) linear equations have no squared variables (x's)
y = (x + 3)(x – 6) – 2x2
Is x squared after you multiply and and subtract
y=x^2-3x-18-2x^2
y=-x^2-3x-18
yes x is squared